Two fair coins are tossed. The probability of getting at least one head is:
- 14
- 34
- 1
- 12
Answer
B) 34
1 - P( TT) = 1 - 14 = 34. SLO M-09-C-05 · Application
Mathematics MCQ · More Probability MCQs · Practise
B) 34
1 - P( TT) = 1 - 14 = 34. SLO M-09-C-05 · Application
Mathematics MCQ · More Probability MCQs · Practise